OpenStudy (anonymous):

3x+ 8y=12 2x+2y=3 part a: create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. show the steps to do this part b: show that the equivalent system has the same soulution as the original system of equations

OpenStudy (mathmale):

I'll let you try this: Multiply the 2nd equation by -4. Now add the first and second equations together and solve for x. Then find y.

OpenStudy (anonymous):

which number in the second equation do i multiply by -4

OpenStudy (mathmale):

You have three terms in the second equation; multiply every one of them by -4. Otherwise you'd be changing the relationship given.

OpenStudy (anonymous):

so its -5+0y=0

OpenStudy (mathmale):

\[-4(2x+2y=3) = -8x-8y=-12.\] Please add this whole equation to the other equation, and then solve the result for x.